OpenStudy (anonymous):

owen scored 85 points on a 100 point test and 26 on a 30 point quiz on which did he get a larger fraction of the total points correct ? show your work

OpenStudy (anonymous):

85/100 = 17/20 = 51/60 26/30 = 13/15 = 52/60 He got a larger fraction on the test where he got 26 on a 30 point test :)